# SDP bounds on quantum codes

### v1

@@ 5. A Lovász bound
We now present a relaxation of the hierarchy that leads to a bound involving the Lovász theta number [Section 3.5]. Consider an intermediate level that is indexed by state monomials of the form $\langle e_{i}\rangle e_{i}$ where $e_{i}$ correspond to letters from $\mathscr{E}_{n}$ [see Eq. (39)]. For simplicity, we will use elements from the $n$-qubit error basis $E_{i}\in\mathcal{E}_{n}$ to index the intermediate level instead of their corresponding letters $e_{i}\in\mathscr{E}_{n}$.
@@ 5.1. An intermediate relaxation level
We now present a fixed level of the hierarchy from Eq. (4.5) with different indexing than before. Similarly to $M^{(\ell)}$, consider a moment matrix $\Gamma$ indexed by the symbols ${\langle E^{\dagger}_{a}\rangle}E_{a}$ for all $E_{a}\in\mathcal{E}_{n}$, where ${\langle E_{a}\rangle}$ inherits the properties of an expectation $\operatorname{tr}(E_{a}\varrho)$ for some state $\varrho$,
@@
with $N=4^{n}-1$ and $\Gamma_{a0}=\Gamma_{aa}$, which approximates ${\langle E^{\dagger}_{a}\rangle}{\langle E_{a}\rangle}$. Now note that if $\Gamma$ is a feasible solution of a semidefinite program with objective function and constraints depending only on the real part of $\Gamma$, then $(\Gamma+\Gamma^{T})/2$ will also be feasible with the same objective value. Without loss of generality, we can therefore consider $\Gamma$ to be real, allowing us to impose the constraint
@@
This follows from the fact that when two tensor-products of Pauli matrices $E_{a}$ and $E_{b}$ anti-commute, the expectation value ${\langle E_{a}E_{b}^{\dagger}\rangle}$ over a quantum state $\varrho$ must be imaginary. Consequently, ${\langle E_{a}^{\dagger}\rangle}{\langle E_{b}\rangle}{\langle E_{a}E_{b}^{\dagger}\rangle}$ is imaginary, and vanishes under the operation $\big{(}\Gamma+\Gamma^{T}\big{)}/2$.
@@ 5.2. Structure constraint
When $E_{a}$ and $E_{b}$ commute, the matrix $\Gamma$ satisfies an extra restriction because it approximates a multiplication of three terms: ${\langle E^{\dagger}_{a}\rangle}$, ${\langle E_{b}\rangle}$, and ${\langle E_{a}E^{\dagger}_{b}\rangle}$. For example, by identifying $E^{\dagger}_{a^{\prime}}=\omega_{a,b}E_{a}E^{\dagger}_{b}$ and $E_{b^{\prime}}=E^{\dagger}_{a}$, we obtain $\omega_{a,b}E_{a^{\prime}}E^{\dagger}_{b^{\prime}}=E_{b}$ which leads to
(60) ${\langle E^{\dagger}_{a}\rangle}{\langle E_{b}\rangle}{\langle E_{a}E^{\dagger}_{b}\rangle}={\langle E^{\dagger}_{a^{\prime}}\rangle}{\langle E_{b^{\prime}}\rangle}{\langle E_{a^{\prime}}E^{\dagger}_{b^{\prime}}\rangle}\,.$
@@
$\Gamma_{ab}=\Gamma_{a^{\prime}b^{\prime}}\,,\quad\quad\quad$ $\text{if}\quad\quad E_{a}E^{\dagger}_{b}=E^{\dagger}_{b}E_{a}\quad\text{and}\quad\,$
(61) $(E^{\dagger}_{a},E_{b},\omega_{ab}E_{a}E^{\dagger}_{b})\quad\text{a permutation of}\quad(E^{\dagger}_{a^{\prime}},E_{b^{\prime}},\omega_{a^{\prime}b^{\prime}}E_{a^{\prime}}E^{\dagger}_{b^{\prime}})\,,$
where $(\omega_{ab}E_{a}E^{\dagger}_{b})\in\mathcal{E}_{n}$ and $\omega_{ab}\in\{\pm 1\}$, since when $E_{a}$ and $E^{\dagger}_{b}$ commute $\omega_{ab}$ cannot be imaginary. Note that this already includes the constraint $\Gamma_{a0}=\Gamma_{aa}$.
@@ 5.3. Projector constraints
As was done for the matrix $M^{(\ell)}$ [see Eq. (54)], we can relax the two projector constraints $\operatorname{tr}(\Pi)=K$ and $\operatorname{tr}(\Pi^{2})=K$ through sums over entries of $\Gamma$,
(62) $\sum^{N}_{a=0}\Gamma_{aa}=\frac{2^{n}}{K},\quad\quad\sum^{N}_{b=0}\Gamma_{ab}=\frac{2^{n}}{K}\Gamma_{aa}\,.$
Summing over both $a$ and $b$ leads to $\sum^{N}_{a,\,b=0}\Gamma_{ab}=\frac{4^{n}}{K^{2}}$.
@@ 5.4. Knill-Laflamme conditions
Similar to Eq (43), we can approximate the Knill-Laflamme conditions $KB_{j}(\varrho)=A_{j}(\varrho)$ if $0\leq j<\delta$ using elements of $\Gamma$ together with $KB_{j}(\varrho)\geq A_{j}(\varrho)$ for any other $j$ [see Remark 5],
(63) $\frac{K}{2^{n}}\sum^{j}_{i=0}K_{j}(i;n)\!\!\sum_{\begin{subarray}{c}E_{a}\in\mathcal{E}_{n}\\ \operatorname{wt}(E_{a})=i\end{subarray}}\Gamma_{a0}\quad\geq\quad\sum_{\begin{subarray}{c}E_{b}\in\mathcal{E}_{n}\\ \operatorname{wt}(E_{b})=j\end{subarray}}\Gamma_{0b}\quad\text{with equality for}\quad 1<j<\delta\,.$
Note that we did not include the condition $KB_{0}(\varrho)=A_{0}(\varrho)$ because it is already approximated by $\sum^{N}_{a=0}\Gamma_{aa}=\frac{2^{n}}{K}$. This can be seen by Eq. (16) that gives $K_{0}(i,n)=1$ for all $i$. Therefore, $K2^{-n}\sum^{n}_{i=0}A_{i}(\varrho)=A_{0}(\varrho)$. Since $A_{0}(\varrho)=1$ and $\Gamma_{a0}=\Gamma_{aa}$, the trace of $\Gamma$ approximates $\sum^{n}_{j=0}A_{j}(\varrho)=2^{n}/K$.
@@ 5.5. Stabilizer codes
Let $S$ be the stabilizer group of a stabilizer code. Recall that $\Gamma$ approximates a matrix with entries $\operatorname{tr}(E_{a}^{\dagger}\varrho)\operatorname{tr}(E_{b}\varrho)\operatorname{tr}(E_{a}E^{\dagger}_{b}\varrho)$. If $\omega_{a}E_{a},\omega_{b}E_{b}\in S$ with $\omega_{a},\omega_{b}=\pm 1$, then $\Gamma_{ab}=1$ since all phases cancel. If $\omega_{a}E\notin S$ or $\omega_{b}E_{b}\notin S$, then $\Gamma_{ab}=0$. Therefore, $\Gamma_{ab}\in\{0,1\}$ for all $a,b\in{\boldsymbol{4}}^{n}$. For a positive semidefinite matrix, the determinant of every $2\times 2$ minor is non-negative, leading to $\Gamma_{ab}\leq\sqrt{\Gamma_{aa}\Gamma_{bb}}$. Due to $\Gamma_{ab}\in\{0,1\}$ this implies that $\Gamma_{ab}\leq\sqrt{\Gamma_{aa}}$. This leads to the following constraint only valid for stabilizer codes
(64) $0\leq\Gamma_{ab}\leq\Gamma_{aa}$
@@ 5.6. Pure codes
In the special case of pure codes, an extra constraint can be added. A code is said to be pure, if the Knill-Laflamme conditions stated in Eq. (8) are strengthened to
(65) $\Pi E^{\dagger}_{a}E_{b}\Pi=0\quad\quad\text{if}\quad\quad 0<\operatorname{wt}(E^{\dagger}_{a}E_{b})<\delta\,.$
As a consequence, we see that $\operatorname{tr}(E^{\dagger}_{a}E_{b}\Pi)=0$ holds for all $0<\operatorname{wt}(E^{\dagger}_{a}E_{b})<\delta$. Because $\Gamma_{ab}={\langle E^{\dagger}_{a}\rangle}{\langle E_{b}\rangle}{\langle E_{a}E_{b}^{\dagger}\rangle}$, this yields the additional constraint,
@@
In particular, choosing $E_{0}=\mathds{1}$ implies that $\Gamma_{aa}=\Gamma_{a0}=0$ for all $E_{a}$ with $0<\operatorname{wt}(E_{a})<\delta$. Eq. (66) strengthens Eq. (63).
@@ 5.7. Main semidefinite program relaxation for qubits
Using the indexing of Eq. (56), we can write the following semidefinite programming relaxation to Eq. (2.1):
@@
$\Gamma_{ab}=\Gamma_{a^{\prime}b^{\prime}}\,,\quad\text{if}\quad E_{a}E^{\dagger}_{b}=E^{\dagger}_{b}E_{a}\quad\text{and}\quad(E^{\dagger}_{a},E_{b},\omega_{ab}E_{a}E^{\dagger}_{b})\quad\,$
a permutation $\hskip 93.89418pt\text{a permutation of}\,\,(E^{\dagger}_{a^{\prime}},E_{b^{\prime}},\omega_{a^{\prime}b^{\prime}}E_{a^{\prime}}E^{\dagger}_{b^{\prime}})\,,$
@@
    (68) $2^{-n}\sum^{n}_{i=0}(-1)^{i}K_{j}(i;n)\sum_{\begin{subarray}{c}E_{a}\in\mathcal{E}_{n}\\ \operatorname{wt}(E_{a})=j\end{subarray}}\Gamma_{aa}\quad\geq\quad 0$
    with equality for $K=1$ and $n-j$ odd [57, Theorem 12].
@@
    For stabilizer codes, we additionally add $0\leq\Gamma_{ab}\leq\Gamma_{aa}$ as shown in Eq. (64) and
@@
    For pure codes, Eq. (66) shows that
@@ 5.8. Lovász bound for self-dual codes
A quantum code with $K=1$ is called self-dual [^3] This is an extension of the nomenclature for stabilizer codes, where $k=0$ ($K=1$) implies the stabilizer group is isomorphic to a self-dual additive classical code.. By definition, such code is pure and we can apply the conditions of the previous section.
@@
Let $G$ be a graph whose vertices correspond to elements of the $n$-qubit Pauli basis $\mathcal{E}_{n}$. Let two vertices $a$ and $b$ be connected, denoted by $a\sim b$, if $E_{a},E_{b}\in\mathcal{E}_{n}$ satisfy $E_{a}E^{\dagger}_{b}+E_{b}E^{\dagger}_{a}=0$ or $0<\operatorname{wt}(E_{a}E^{\dagger}_{b})<\delta$; and let $a\sim a$ if $0<\operatorname{wt}(E_{a})<\delta$.
This simply incorporates the two constraints of Eq. (59) and (66). To see this, note that if a positive semidefinite matrix has an element in the diagonal equal to zero, this implies that all elements in the same row and column are zero, too. With this notation, we relax the SDP from Eq. (5.7) for self-dual codes to:
@@
Because of $a\sim a$ if and only if $a\sim 0$ the last constraint can be restricted to $a,b\neq 0$ without changing the objective value. Thus SDP (5.8) is nothing else than the semidefinite program for the Lovász theta number (3.5) of the confusability graph $G$ (Def. 8) with the vertex corresponding to the identity removed [^4] In fact, this is an instance of a Lovász theta number for a hypergraph [10].. To see this, note that $\Gamma$ is indexed by $0,\dots,N$, while the sum in Eq. (5.8) starts at $1$, corresponding to the sum over $N$ elements in Eq. (3.5). For a self-dual code to exist, the constraint $\sum^{N}_{a=0}\Gamma_{aa}=\frac{2^{n}}{K}$ in Eq. (5.7) must be met. Thus the objective value of Eq. (5.8) must be larger or equal than $12^{n}-$ 1. As a consequence, we have the following corollary:
@@
If a pure $(\!(n,1,\delta)\!)_{2}$ code exists, then $2^{n}\leq 1+\vartheta(G^{\prime})$, where $G^{\prime}$ is the confusability graph $G$ (c.f. Def. 8) of a self-dual quantum code with the identity vertex removed.
@@
For the confusability graph $G^{\prime}$ of the code $(\!(4,1,3)\!)_{2}$, the Lovász number is $\vartheta(G^{\prime})=7$. By Corollary 9, $1+\vartheta(G^{\prime})<2^{4}$ implies that a $(\!(4,1,3)\!)_{2}$ code does not exist. While this result was already known from the linear programming bounds that use the shadow inequalities [see Eq. (3.4)] 555 The weights $(1,0,0,12,3)$ are non-negative and invariant under the quantum MacWillams transformation., we think that Corollary 9 provides a conceptually simpler approach.
Note that $\Gamma$ from Eq. (5.8) has zeros in all rows and columns with $0<\operatorname{wt}(E_{a})<\delta$. The positivity of $\Gamma$ is then equivalent to the positivity of the same matrix without such rows and columns. This allows us to state the following.
@@
Let $G^{\prime\prime}$ be the subgraph of the confusability graph $G$ from Def. 8, whose vertices $\mathds{1}$ and $E_{a}\in\mathcal{E}_{n}$ with $\operatorname{wt}(E_{a})<\delta$ are removed. Then, $\vartheta(G^{\prime\prime})=\vartheta(G^{\prime})$.

### v2

@@ 5. A Lovász bound
We now present a relaxation of the hierarchy from Eq. (4.5) that leads to a bound involving the Lovász theta number [Section 3.5]. The relaxation is indexed by state monomials of the form $\langle e_{a}\rangle e_{a}$ where the $e_{a}$ correspond to letters from $\mathscr{E}_{n}$ [see Eq. (40)]. For readability, we will use elements from the $n$-qubit Pauli basis $E_{a}\in\mathcal{E}_{n}$ for indexing rather the corresponding letters $e_{a}\in\mathscr{E}_{n}$.
@@ 5.1. An intermediate relaxation level
Consider a moment matrix $\Gamma$ indexed by the symbols ${\langle E^{\dagger}_{a}\rangle}E_{a}$ with $E_{a}\in\mathcal{E}_{n}$, where $\langle E_{a}\rangle$} inherits the properties of an expectation $\operatorname{tr}(E_{a}\varrho)$ for some state $\varrho$,
@@
with $N=4^{n}-1$ and $\Gamma_{a0}=\Gamma_{aa}$, which represents $\operatorname{tr}(E^{\dagger}_{a}\varrho)\operatorname{tr}(E_{a}\varrho)$. Now note that if $\Gamma$ is a feasible solution of a semidefinite program with objective function and constraints that depend only on the real part of $\Gamma$, then $\operatorname{Re}(\Gamma)=(\Gamma+\Gamma^{T})/2$ is also feasible with the same objective value. Without loss of generality, we can therefore restrict $\Gamma$ to be real. This leads to the constraint
@@
This follows from the fact that when two Pauli strings $E_{a}$ and $E_{b}$ anti-commute, then $E_{a}E_{b}^{\dagger}$ is imaginary.
@@ 5.2. Structure constraint
When $E_{a}$ and $E_{b}$ commute, the matrix $\Gamma$ satisfies an additional constraint related to a multiplication of three terms: $\operatorname{tr}(E^{\dagger}_{a}\varrho)$, $\operatorname{tr}(E_{b}\varrho)$, and $\operatorname{tr}(E_{a}E^{\dagger}_{b}\varrho)$. For example, if $E^{\dagger}_{c^{\prime}}=\omega_{ab}E_{a}E^{\dagger}_{b}$ and $E_{d^{\prime}}=E^{\dagger}_{a}$, then $\omega_{ab}E_{c^{\prime}}E^{\dagger}_{d^{\prime}}=E_{b}$, leading to the condition
(61) ${\langle E^{\dagger}_{a}\rangle}{\langle E_{b}\rangle}{\langle E_{a}E^{\dagger}_{b}\rangle}={\langle E^{\dagger}_{c^{\prime}}\rangle}{\langle E_{d^{\prime}}\rangle}{\langle E_{c^{\prime}}E^{\dagger}_{d^{\prime}}\rangle}\,.$
@@
$\Gamma_{ab}=\Gamma_{c^{\prime}d^{\prime}}\,$$\quad\quad\quad$ $\text{if}\quad\quad E_{a}E^{\dagger}_{b}-E^{\dagger}_{b}E_{a}=0\quad\text{and}\quad\,$
(62) $(E^{\dagger}_{a},E_{b},\omega_{ab}E_{a}E^{\dagger}_{b})\quad\text{a permutation of}\quad(E^{\dagger}_{c^{\prime}},E_{d^{\prime}},\omega_{c^{\prime}d^{\prime}}E_{c^{\prime}}E^{\dagger}_{d^{\prime}})\,$,
where $(\omega_{ab}E_{a}E^{\dagger}_{b})\in\mathcal{E}_{n}$ and $\omega_{ab}\in\{\pm 1\}$, since when $E_{a}$ and $E^{\dagger}_{b}$ commute $\omega_{ab}$ must be real. Note that Eq. (5.2) includes $\Gamma_{a0}=\Gamma_{aa}$ as a special case.
@@ 5.3. Projector constraints
As done for the matrix $M^{(\ell)}$ in Eq. (51), we relax the projector constraints $\operatorname{tr}(\rho^{2})=1/K$ and $\operatorname{tr}(\rho^{3})=1/K^{2}$ by expressing them, respectively, through sums over elements of $\Gamma$,
(63) $\sum^{N}_{a=0}\Gamma_{aa}=\frac{2^{n}}{K}\,,\quad\quad\quad\sum^{N}_{a,b=0}\Gamma_{ab}=\frac{4^{n}}{K^{2}}\,.$
Another constraint follows from $\operatorname{tr}(E_{a}\varrho^{2})=\frac{1}{K}\operatorname{tr}(E_{a}\varrho)$ as shown in Remark 6: the sum $\sum^{N}_{b=0}\Gamma_{ab}$ approximates
$\sum^{N}_{b=0}\operatorname{tr}(E^{\dagger}_{a}\rho)\operatorname{tr}(E_{b}\rho)\operatorname{tr}(E_{a}E^{\dagger}_{b}\rho)$ $=\operatorname{tr}(\rho E^{\dagger}_{a})\sum^{N}_{b=0}\operatorname{tr}((E_{b}\otimes E_{a}E^{\dagger}_{b})\rho^{\otimes 2})$
$=\operatorname{tr}(\rho E^{\dagger}_{a})\operatorname{tr}((\mathds{1}\otimes E_{a})\sum^{N}_{b=0}(E_{b}\otimes E^{\dagger}_{b})\rho^{\otimes 2})$
$=\operatorname{tr}(\rho E^{\dagger}_{a})\operatorname{tr}((\varrho\otimes\varrho E_{a})(1,2))$
(64) $=2^{n}\operatorname{tr}(\rho E^{\dagger}_{a})\operatorname{tr}(\rho^{2}E_{a})\,$,
where we used that $\sum_{E_{a}\in\mathcal{E}_{n}}E_{a}\otimes E^{\dagger}_{a}=2^{n}(1,2)$. Since $\operatorname{tr}(E_{a}\varrho^{2})=(1/K)\operatorname{tr}(E_{a}\varrho)$, we obtain
(65) $\sum^{N}_{b=0}\Gamma_{ab}=\frac{2^{n}}{K}\Gamma_{aa}\,.$
Note that Eq. (65) implies the first and second conditions of Eq. (63), which corresponds to the projector constraint of degree two, $\operatorname{tr}(\rho^{2})=1/K$, and three, $\operatorname{tr}(\rho^{3})=1/K^{2}$, respectively. The first is obtained by setting $a=0$ in the Eq. (65) and noting that $\Gamma_{aa}=\Gamma_{a0}$ and $\Gamma_{00}=1$; the second is obtained by summing Eq. (65) over all values of $a$ and taking into count the first condition of Eq. (63).
@@ 5.4. Knill-Laflamme conditions
Similar to Eq. (44), we can approximate the Knill-Laflamme conditions $KB_{j}(\varrho)=A_{j}(\varrho)$ if $0\leq j<\delta$ using elements of $\Gamma$ together with $KB_{j}(\varrho)\geq A_{j}(\varrho)$ for $j>\delta$ [see Remark 5],
(66) $\frac{K}{2^{n}}\sum^{n}_{i=0}K_{j}(i;n)\!\!\sum_{\begin{subarray}{c}E_{a}\in\mathcal{E}_{n}\\ \operatorname{wt}(E_{a})=i\end{subarray}}\Gamma_{a0}\geq\sum_{\begin{subarray}{c}E_{b}\in\mathcal{E}_{n}\\ \operatorname{wt}(E_{b})=j\end{subarray}}\Gamma_{0b}\quad\quad\text{with equality for}\quad 1<j<\delta\,.$
Note that we did not include the condition $KB_{0}(\varrho)=A_{0}(\varrho)$ because it is already approximated by $\sum^{N}_{a=0}\Gamma_{aa}=2^{n}/K$. This can be seen using Eq. (16), which gives $K_{0}(i,n)=1$ for all $i$. Therefore, $K2^{-n}\sum^{n}_{i=0}A_{i}(\varrho)=A_{0}(\varrho)$. Since $A_{0}(\varrho)=1$ and $\Gamma_{a0}=\Gamma_{aa}$, the trace of $\Gamma$ approximates $\sum^{n}_{j=0}A_{j}(\varrho)=2^{n}/K$.
@@ 5.5. Stabilizer codes
Let $S$ be the stabilizer group of a stabilizer code. For a stabilizer code, $\Gamma$ represents $\chi\chi^{T}$ where $\chi$ is the incidence vector of the stabilizer group. This can be seen in the following way: given a code, $\Gamma$ has entries $\operatorname{tr}(E_{a}^{\dagger}\varrho)\operatorname{tr}(E_{b}\varrho)\operatorname{tr}(E_{a}E^{\dagger}_{b}\varrho)$. But $\operatorname{tr}(E_{a}\varrho)=1$ if $E_{a}\in S$, and $\operatorname{tr}(E_{a}\varrho)=0$ otherwise. Thus $\Gamma_{ab}=1$ if both $E_{a},E_{b}\in S$ and zero otherwise.
We can use this fact in the following way. For a positive semidefinite matrix, the determinant of every $2\times 2$ minor is non-negative, leading to $\Gamma_{ab}\leq\sqrt{\Gamma_{aa}\Gamma_{bb}}$. Due to $\Gamma_{ab}\in\{0,1\}$ this implies that $\Gamma_{ab}\leq\sqrt{\Gamma_{aa}}$, giving a constraint for stabilizer codes,
(67) $0\leq\Gamma_{ab}\leq\Gamma_{aa}\,.$
@@ 5.6. Pure codes
In the special case of pure codes, an extra constraint can be added. Recall that a code is said to be pure if $\operatorname{tr}(E^{\dagger}_{a}E_{b}\Pi)=0$ for all $0<\operatorname{wt}(E^{\dagger}_{a}E_{b})<\delta$. Because $\Gamma_{ab}={\langle E^{\dagger}_{a}\rangle}{\langle E_{b}\rangle}{\langle E_{a}E_{b}^{\dagger}\rangle}$, we have additionally that
@@
In particular, choosing $b=0$ (since $E_{0}=\mathds{1}$) implies that $\Gamma_{aa}=\Gamma_{a0}=0$ for all $a$ with $0<\operatorname{wt}(E_{a})<\delta$. Eq. (68) strengthens Eq. (66).
@@ 5.7. Main semidefinite program relaxation for qubits
Let $\Gamma$ be indexed as in Eq. (57) and without loss of generality let $\Gamma$ be real. Consider a SDP relaxation to Eq. (2.1):
@@
$\Gamma_{ab}=\Gamma_{cd}\,\,\quad\text{if}\quad E_{a}E^{\dagger}_{b}=E^{\dagger}_{b}E_{a}\quad\text{and}\quad(E^{\dagger}_{a},E_{b},\omega_{ab}E_{a}E^{\dagger}_{b})\quad\,$
$\hskip 93.89418pt\text{a permutation of}\,\,(E^{\dagger}_{c},E_{d},\omega_{cd}E_{c}E^{\dagger}_{d})\,$,
@@
    (70) $2^{-n}\sum^{n}_{i=0}(-1)^{i}K_{j}(i;n)\sum_{\begin{subarray}{c}E_{a}\in\mathcal{E}_{n}\\ \operatorname{wt}(E_{a})=i\end{subarray}}\Gamma_{aa}\quad\geq\quad 0$
    with equality for $K=1$ and njn-j odd [62, Theorem 12].
@@
    For stabilizer codes add $0\leq\Gamma_{ab}\leq\Gamma_{aa}$ [Eq. (67)] and [9, 70]
@@
    For pure codes add [Eq. (68)],
@@ 5.8. Lovász bound for self-dual codes
Recall that a quantum code with $K=1$ is called self-dual. By definition, such code is pure and we can apply the conditions of the previous section.
@@
Let $G$ be a graph whose vertices correspond to elements of the $n$-qubit Pauli basis $\mathcal{E}_{n}$. Let two vertices $a$ and $b$ be connected, written $a\sim b$, if $E_{a},E_{b}\in\mathcal{E}_{n}$ satisfy $E_{a}E^{\dagger}_{b}+E_{b}E^{\dagger}_{a}=0$ or $0<\operatorname{wt}(E_{a}E^{\dagger}_{b})<\delta$; and let $a\sim a$ if $0<\operatorname{wt}(E_{a})<\delta$.
This simply incorporates the two constraints of Eq. (60) and (68). To see this, note that if a positive semidefinite matrix has an element in the diagonal equal to zero, then all elements in the same row and column are zero, too. With this notation, we relax the SDP from Eq. (5.7) [that include the pure code conditions in Eq. (72)] for self-dual codes to:
@@
Because of $a\sim a$ if and only if $a\sim 0$ the last constraint can be restricted to $a,b\neq 0$ without changing the objective value. Thus the SDP in Eq. (5.8) is nothing other than the semidefinite program for the Lovász theta number [Eq. (3.5)] of the confusability graph $G$ with the vertex corresponding to the identity removed. To see this, note that $\Gamma$ is indexed by $0,\dots,N$, while the sum in Eq. (5.8) starts at $1$, corresponding to the sum over $N$ elements in Eq. (3.5). For a self-dual code to exist, the constraint $\sum^{N}_{a=0}\Gamma_{aa}={2^{n}/K}$ in Eq. (5.7) must be met. Thus the objective value of Eq. (5.8) must be larger or equal than $2^{n}-1$. As a consequence, we have the following corollary:
@@
If a pure $(\!(n,1,\delta)\!)_{2}$ code exists, then $2^{n}\leq 1+\vartheta(G^{\prime})$, where $G^{\prime}$ is the confusability graph $G$ [Definition 9] of a self-dual quantum code with the identity vertex removed.
@@
For the confusability graph $G^{\prime}$ of a $(\!(4,1,3)\!)_{2}$ code, the Lovász theta number is $\vartheta(G^{\prime})=7$. By Corollary 10, $1+\vartheta(G^{\prime})<2^{4}$ implies that a $(\!(4,1,3)\!)_{2}$ code does not exist. While this result was already known from the linear programming bounds enhanced with the shadow inequalities [Eq. (3.4)] 101010 The weights $[1,0,0,12,3]$ are unique in being non-negative, invariant under the quantum MacWillams transformation, and satisfying the pure code constraint. However, they violate the shadow inequalities., we think that using Corollary 10 provides a conceptually simpler approach.
Note that $\Gamma$ from Eq. (5.8) has zeros in all rows and columns with $0<\operatorname{wt}(E_{a})<\delta$. The positivity of $\Gamma$ is then equivalent to the positivity of the same matrix without such rows and columns:
@@
Let $G^{\prime\prime}$ be the subgraph of the confusability graph $G$ from Definition 9, whose vertices $\mathds{1}$ and $E_{a}\in\mathcal{E}_{n}$ such that $\operatorname{wt}(E_{a})<\delta$ removed. Then, $\vartheta(G^{\prime\prime})=\vartheta(G^{\prime})$.